OpenStudy (anonymous):

Solve the equation. x^2 = 150

OpenStudy (studygurl14):

Do the opposite of squaring to both sides of the equation.

OpenStudy (anonymous):

let me get the answers

OpenStudy (studygurl14):

That would be taking the square root, so... \(\Large x^2=150\rightarrow \sqrt{x^2}=\sqrt{150}\rightarrow x = \sqrt{150}\rightarrow ?\) You can simplify the \(\sqrt{150}\) more.

OpenStudy (jackofalltradez):

Square root each side: